# Irrationality of rapidly converging series: a problem of Erdős and Graham

### v1

@@ Abstract.
Answering a question of Erdős and Graham, we show that the double exponential growth condition $\limsup_{n\to\infty}a_{n}^{1/\phi^{n}}=\infty$ for a monotonically increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$, together with the bound $a_{n}a_{n+1}\geq n^{1+\tau}$, is sufficient for the series $\sum_{n=1}^{\infty}1/(a_{n}a_{n+1})$ to have an irrational sum; here $\phi$ denotes the golden ratio $(1+\sqrt{5})/2$ and $\tau>0$. We also provide a positive generalization to $\sum_{n=1}^{\infty}1/(a_{n}^{w_{0}}\cdots a_{n+d-1}^{w_{d-1}})$, and a negative result showing that some of its instances are essentially optimal. The original problem was autonomously solved by the AI agent Aletheia, powered by Gemini Deep Think, while the remaining material is largely a product of human-AI interactions.
@@ 1. Declaration of AI usage
    Aletheia, a custom math-research agent built on top of Gemini Deep Think (whose development was led by Tony Feng), found a solution to Question 1. From the original output, it was clear that some generalization was possible. Specifically, the proof used rough estimates such as $1.7<2$, $\log x<x$, etc.
@@ 2. Introduction
The present paper provides a positive answer to Question 1 and its natural generalizations, namely Theorems 2 and 3. The proposers of Question 1, Erdős and Graham, noted that hypothesis (2.1) “should be enough” to imply the conclusion, while remarking that “it is not too hard to show that if $a_{n}\to\infty$ rapidly enough” [4, p. 64]. Their latter remark can be inferred from [8, Theorem 1] to mean, for example, that $\lim_{n\to\infty}a_{n}^{1/2^{n}}=\infty$. They also asked “what the strongest theorem of this type” would be and our Theorem 2 below provides a satisfactory answer.
@@
for some positive integer sequences $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$ with $a_{n}\geq 2$ for every $n$ are known as Cantor series. Investigation of their irrationality dates back to the mid-$19$th century [2, 20, 6, 7], while more recent relevant literature is [21, 15, 14]. These results can solve particular cases of Question 1 simply by rewriting (2.2) as the Cantor series (2.3) with $b_{n}=a_{1}\cdots a_{n-2}$. In this way, a result by Hančl and Tijdeman [15, Theorem 3.2] confirms that (2.2) is not a rational number as soon as we make the following assumptions on divisibility and the “relative growth” among the terms $a_{n}$:
@@
    If a monotonically increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ satisfies
@@
We believe that Theorem 2 is a novel result when $d\geq 2$. For $d=1$ we get $\psi=2$, so Part (1) is an easy consequence of Erdős’s theorem [8, Theorem 1], while the well-known Sylvester sequence provides an explicit example of a sequence as in Part (2); see the remarks in [17, Section 1]. For $d\geq 2$, Part (1) is a special case of an even more general Theorem 3 below, while Part (2) is a particular instance of Theorem 5.
@@ 3. Proof of Theorem 3
    If $x_{k}\geq k^{1+\tau}$ for all $k\in[Q,R]$ with some $\tau>\eta$, then _k=Q^R ykxk1xQ(τ-η)/1+τ.
@@ 4. Proof of Theorem 5
Lemma 14 does not seem to appear in the literature, but somewhat similar constructions have been used in the context of irrationality of Ahmes and Cantor series, for instance in relation to [16, Theorems 1–3][21, Example 5.2][15, Algorithm 3.1], and [17, Lemma 5.1].
@@ Acknowledgements
The third named author is supported by Mid-Career Researcher Program (RS-2023-00278510) through the National Research Foundation funded by the government of Korea, and also by KIAS Individual Grants (MG073601 and MG084001, respectively) at Korea Institute for Advanced Study. The fourth named author is supported in part by the Croatian Science Foundation under the project HRZZ-IP-2022-10-5116 (FANAP). We thank Thomas Bloom for creating and maintaining the website Erdős problems and Terence Tao for valuable comments on an earlier version of this paper. We are thankful to the authors of [10] for building the Aletheia agent and letting the authors use it.

### v2

@@ Abstract.
Answering a question of Erdős and Graham, we show that the double exponential growth condition $\limsup_{n\to\infty}a_{n}^{1/\phi^{n}}=\infty$ for a strictly increasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ is sufficient for the series $\sum_{n=1}^{\infty}1/(a_{n}a_{n+1})$ to have an irrational sum; here $\phi$ denotes the golden ratio. We also provide a positive generalization to $\sum_{n=1}^{\infty}1/(a_{n}^{w_{0}}\cdots a_{n+d-1}^{w_{d-1}})$, and a negative result showing that some of its instances are essentially optimal. The original problem was autonomously solved by the AI agent Aletheia, powered by Gemini Deep Think, while the remaining material is largely a product of human-AI interactions.
@@ 1. Declaration of AI usage
    Aletheia [10] [^1] Aletheia is a custom mathematics research agent built upon Gemini Deep Think, details at https://github.com/google-deepmind/superhuman/tree/main/aletheia. found a solution to Question 1. From the original output, it was clear that some generalization was possible. Specifically, the proof used rough estimates such as $1.7<2$, $\log x<x$, etc. See [11] for relevant raw outputs.
@@ 2. Introduction
The present paper provides a positive answer to Question 1 and its natural generalizations, namely Theorems 2 and 3. The proposers of Question 1, Erdős and Graham, noted that hypothesis (2.1) “should be enough” to imply the conclusion, while remarking that “it is not too hard to show that if $a_{n}\to\infty$ rapidly enough” [4, p. 64]. Their latter remark can be inferred from [8, Theorem 1] to mean, for example, that $\lim_{n\to\infty}a_{n}^{1/2^{n}}=\infty$. They also asked “what the strongest theorem of this type” would be, and our result, Theorem 2 below, provides a satisfactory answer.
@@
for some positive integer sequences $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$ with $a_{n}\geq 2$ for every $n$ are known as Cantor series. Investigation of their irrationality dates back to the mid-19th century [2, 21, 6, 7], while more recent relevant literature is [22, 16, 15]. These results can solve particular cases of Question 1 simply by rewriting (2.2) as the Cantor series (2.3) with $b_{n}=a_{1}\cdots a_{n-2}$. In this way, a result by Hančl and Tijdeman [16, Theorem 3.2] confirms that (2.2) is not a rational number as soon as we make the following assumptions on divisibility and the “relative growth” among the terms $a_{n}$:
@@
    If a monotonically increasing 222i.e, non-decreasing sequence of positive integers $\{a_{n}\}_{n=1}^{\infty}$ satisfies
@@
We believe that Theorem 2 is a novel result when $d\geq 2$. For $d=1$ we get $\psi=2$, so Part (1) is an easy consequence of Erdős’s theorem [8, Theorem 1], while the well-known Sylvester sequence provides an explicit example of a sequence as in Part (2); see the remarks in [18, Section 1]. For $d\geq 2$, Part (1) is a special case of an even more general Theorem 3 below, while Part (2) is a particular instance of Theorem 5.
@@
  3. (3)
    Another consequence of Theorem 3 is the following. Suppose $d\geq 2$ and $\psi^{d}=\psi^{d-1}+1$ for some $\psi>0$. If $\{a_{n}\}_{n=1}^{\infty}$ is a strictly increasing sequence of positive integers satisfying lim sup_n→∞ a_n^1/ψ^n=∞, then the sum ∑_n=1^∞1anan+1⋯an+d-1 is irrational. To see this, note first that $a_{n}\geq n$, and hence, a_na_n+1⋯a_n+d-1≥n^d≥n^2. So, the hypotheses of Theorem 3 are automatically satisfied. In particular, this proves the result mentioned in the abstract.
@@ 3. Proof of Theorem 3
    If $x_{k}\geq k^{1+\tau}$ for all $k\in[Q,R]$ with some $\tau>\eta$, then _k=Q^R ykxk1xQ(τ-η)/(1+τ).
@@ 4. Proof of Theorem 5
Lemma 14 does not seem to appear in the literature, but somewhat similar constructions have been used in the context of irrationality of Ahmes and Cantor series, for instance in relation to [17, Theorems 1–3][22, Example 5.2][16, Algorithm 3.1], and [18, Lemma 5.1].
@@ Acknowledgements
The third named author is supported by Mid-Career Researcher Program (RS-2023-00278510) through the National Research Foundation funded by the government of Korea, and also by KIAS Individual Grants (MG073601 and MG084001, respectively) at Korea Institute for Advanced Study. The fourth named author is supported in part by the Croatian Science Foundation under the project HRZZ-IP-2022-10-5116 (FANAP). We thank Thomas Bloom for creating and maintaining the website Erdős problems and Terence Tao for valuable comments on an earlier draft of this paper. We thank the authors of [10] for building the Aletheia agent and for allowing the authors of this work to utilize it.